OpenStudy (anonymous):

solve by the substitution method 6x+9y=6 -5x+y=29

OpenStudy (a_clan):

put y=29+5x in first equation

OpenStudy (anonymous):

6x+9(5x+29)=6 6x+45x+261=6 51x=-255 x=-5 y=29+5x substitute x=-5 y=29-25=4 x=-5 and y=4